OpenStudy (anonymous):

14z^2/sqrt(z^9) = 14 sqrt(z)/z^7 , right?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Use the law of indices. sqrt a = a^0.5 a^m / a^n = a^(m - n)

OpenStudy (anonymous):

I don't use indices in my class.. I don't even know what that is...

OpenStudy (anonymous):

You are rationalising? It should be 14 sqrt (z^9) / z^7

OpenStudy (anonymous):

No, finding the exact value....

jimthompson5910 (jim_thompson5910):

do you know how to simplify \[\large \sqrt{z^9}\] ??

OpenStudy (anonymous):

I was told to subtract the exponents. :P I don't know how to simplify them

jimthompson5910 (jim_thompson5910):

that step will come next, but first we must simplify \[\large \sqrt{z^9}\]

jimthompson5910 (jim_thompson5910):

doing that will give us this \[\large \sqrt{z^9}\] \[\large \sqrt{z^8*z}\] \[\large \sqrt{z^8}*\sqrt{z}\] \[\large (z^8)^{1/2}*\sqrt{z}\] \[\large z^{8*1/2}*\sqrt{z}\] \[\large z^{8/2}*\sqrt{z}\] \[\large z^4\sqrt{z}\]

jimthompson5910 (jim_thompson5910):

so \[\large \sqrt{z^9}\] turns into \[\large z^4\sqrt{z}\] hopefully you can see how

OpenStudy (anonymous):

Yes, kind of, thank you.

jimthompson5910 (jim_thompson5910):

so this means that \[\large \frac{14z^2}{\sqrt{z^9}\] turns into \[\large \frac{14z^2}{z^4\sqrt{z}\] now simplify and then rationalize the denominator